# vMFCoOp: Towards Equilibrium on a Unified Hyperspherical Manifold for Prompting Biomedical VLMs

### v1

@@ Abstract
Recent advances in context optimization (CoOp) guided by large language model (LLM)–distilled medical semantic priors offer a scalable alternative to manual prompt engineering and full fine-tuning for adapting biomedical CLIP-based vision-language models (VLMs). However, prompt learning in this context is challenged by semantic misalignment between LLMs and CLIP variants due to divergent training corpora and model architectures; it further lacks scalability across continuously evolving families of foundation models. More critically, pairwise multimodal alignment via conventional Euclidean-space optimization lacks the capacity to model unified representations or apply localized geometric constraints, which tends to amplify modality gaps in complex biomedical imaging and destabilize few-shot adaptation. In this work, we propose vMFCoOp, a framework that inversely estimates von Mises–Fisher (vMF) distributions on a shared Hyperspherical Manifold, aligning semantic biases between arbitrary LLMs and CLIP backbones via Unified Semantic Anchors to achieve robust biomedical prompting and superior few-shot classification. Grounded in three complementary constraints, vMFCoOp demonstrates consistent improvements across 14 medical datasets, 12 medical imaging modalities, and 13 anatomical regions, outperforming state-of-the-art methods in accuracy, generalization, and clinical applicability. This work will be continuously expanded to encompass more downstream applications, and the corresponding resources are intended to be shared through https://github.com/VinyehShaw/UniEqui.
@@ Introduction
Our vMFCoOp framework flexibly plugs into any LLM and biomedical-CLIP backbone to jointly learn prompt contexts. It imposes complementary constraints on a shared hyperspherical manifold, guiding prompt and visual embeddings toward a unified semantic anchor field reconciling multimodal textual biases via the vMF Inverse Mapping Estimation, in turn mitigating textual semantic transfer misalignment across diverse model priors and granularity levels in biomedical imaging. Figure 1B provides an intuitive lower-dimensional overview of the hyperspherical manifold, highlighting three hierarchical constraints: Semantic Anchor Loss ($\mathcal{L}_{\mathit{\boldsymbol{anc}}}$) encourages the prompt context embeddings to converge toward a unified semantic anchor by calibrating the inherently mismatched LLM and CLIP embedding spaces into an aligned subspace, Spherical Contrastive Loss ($\mathcal{L}_{\mathit{\boldsymbol{sc}}}$) introduces supervised angular margins over the unit hypersphere to enhance inter-class separability within the shared manifold, and Symmetric Cross-Entropy Loss ($\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$) jointly minimizes forward and reverse divergences between image- and prompt-based class distributions. Consequently, few-shot samples converge toward multimodal semantic equilibrium via this unified anchor-guided manifold space to facilitate class-level prompt prediction.
@@
We use BiomedCLIP and other biomedical CLIP variants take a tokenized prompt sequence $\boldsymbol{X_{t}}\in\mathbb{R}^{B\times L}$ and a preprocessed RGB medical image $\boldsymbol{X_{v}}\in\mathbb{R}^{B\times 3\times H\times W}$ as input, and encodes them into the embedding space via a text encoder $\boldsymbol{E_{t}}$ and a vision encoder $\boldsymbol{E_{v}}$, respectively. The text $\boldsymbol{T}$ and vision embeddings $\boldsymbol{V}$ are given by $\boldsymbol{E_{t}(X_{t})}\in\mathbb{R}^{C\times D}$ and $\boldsymbol{E_{v}(X_{v})}\in\mathbb{R}^{B\times D}$, where $B$ is the batch size, $L$ is the maximum token length, $H$ and $W$ are image dimensions, $C$ is the number of classes, and $D$ is the embedding dimension. Our prompt learner constructs class‑specific prompts of the form “a photo of a[CLASS].”, where “a photo of a” serves as the initialized learnable context, and [CLASS] denotes the class name token. Classification is performed by computing cosine similarities $\cos(,)$ between the image features $V$ and each class prompt embedding $\boldsymbol{T_{p}^{(i)}}$, scaling these scores by the temperature $\tau$, applying $\mathrm{Softmax}\bigl(\cos(\boldsymbol{V},\boldsymbol{T_{p}^{(i)}})/\tau\bigr)$ to obtain class probabilities $p_{i}$, and selecting the final label as $\arg\max_{i}\,p(y=i\mid\boldsymbol{V,T_{p}^{(i)}})$. More mathematical derivations and additional theoretical proofs are provided in the Proofs document.
Following the setup of BiomedCoOp, we generate prompts for each class $c_{i}$ in medical modality by querying each of several recent LLMs separately with the standardized instruction:“Give N textual descriptions of visual discriminative features for distinct medical cases of [CLASS] found in [MODALITY].” We treat all LLM-provided template features as samples distributed on a shared manifold space and estimate a class-level semantic prototype field over this space. For each class $c_{i}$, we extract a semantic prototype vector $\boldsymbol{\mu}_{L,c}^{(i)}$, which guides the prompt learner in balancing multi-modal alignment and semantic expressiveness.
@@ 2. How to avoid distractors? We apply a row-wise softmax cross-entropy across $C$ classes:
Figure 2: Effect of prompting variations on saliency maps, where (a)–(e) illustrate different strategies (zoom in for details). Note that the fourth row depicts a rare cardiac cine MRI case in which the patient has a posterior mediastinal tumor, simulating a few-shot fine-tuning scenario with limited data and reflecting real clinical settings. BiomedCoOp successfully localizes the approximate lesion region under such challenging conditions.
@@
We summarize the three hierarchical constraints that define the core components of vMFCoOp. In this ablation, we systematically isolate each component under identical settings: $\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$ denotes replacing standard cross-entropy with Symmetric Cross-Entropy Loss atop BiomedCLIP; $\mathcal{L}_{\mathit{\boldsymbol{anc}}}$ introduces a unified semantic anchor field via inverse vMF estimation on the hyperspherical manifold; and $\mathcal{L}_{\mathit{\boldsymbol{sc}}}$ adds Spherical Contrastive Loss to enforce angular margins. As shown in Table 2, each constraint provides incremental improvements, while their combination yields the strongest gains. Notably, incorporating $\mathcal{L}_{\mathit{\boldsymbol{anc}}}$ consistently enhances both few-shot performance and base-to-novel generalization, underscoring the importance of aligning multimodal representations via a shared semantic anchor. Adding $\mathcal{L}_{\mathit{\boldsymbol{sc}}}$ further improves novel class accuracy, especially in higher-shot regimes, by reducing overfitting through explicit angular separation. While $\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$ alone offers modest improvements, its integration with the other two constraints amplifies robustness and cross-modal alignment. These results collectively demonstrate that the three constraints are complementary and jointly contribute to optimal generalization and stability in biomedical few-shot learning.
@@ Conclusion
Acknowledgments
This research was conducted using data from the UK Biobank Resource under Application Number 603483, as part of an existing approved project. The authors express sincere gratitude to all participants and research personnel involved in the UK Biobank initiative.

### v3

@@ Abstract
Recent advances in context optimization (CoOp) guided by large language model (LLM)–distilled medical semantic priors offer a scalable alternative to manual prompt engineering and full fine-tuning for adapting biomedical CLIP-based vision-language models (VLMs). However, prompt learning in this context is challenged by semantic misalignment between LLMs and CLIP variants due to divergent training corpora and model architectures; it further lacks scalability across continuously evolving families of foundation models. More critically, pairwise multimodal alignment via conventional Euclidean-space optimization lacks the capacity to model unified representations or apply localized geometric constraints, which tends to amplify modality gaps in complex biomedical imaging and destabilize few-shot adaptation. In this work, we propose vMFCoOp, a framework that inversely estimates von Mises–Fisher (vMF) distributions on a shared Hyperspherical Manifold, aligning semantic biases between arbitrary LLMs and CLIP backbones via Unified Semantic Anchors to achieve robust biomedical prompting and superior few-shot classification. Grounded in three complementary constraints, vMFCoOp demonstrates consistent improvements across 14 medical datasets, 12 medical imaging modalities, and 13 anatomical regions, outperforming state-of-the-art methods in accuracy, generalization, and clinical applicability. This work aims to continuously expand to encompass more downstream applications, and the corresponding resources are intended to be shared through https://github.com/VinyehShaw/UniEqui.
@@ Introduction
Our vMFCoOp framework flexibly plugs into any LLM and biomedical-CLIP backbone to jointly learn prompt contexts. It imposes complementary constraints on a shared hyperspherical manifold, guiding prompt and visual embeddings toward a unified semantic anchor field reconciling multimodal textual biases via the vMF Inverse Mapping Estimation, in turn mitigating textual semantic transfer misalignment across diverse model priors and granularity levels in biomedical imaging. Figure 1B provides an intuitive lower-dimensional overview of the hyperspherical manifold, highlighting three hierarchical constraints: Semantic Anchor Loss ($\mathcal{L}_{\mathit{\boldsymbol{anc}}}$) encourages the prompt context embeddings to converge toward a unified semantic anchor by calibrating the inherently mismatched LLM and CLIP embedding spaces into an aligned subspace, Spherical Contrastive Loss ($\mathcal{L}_{\mathit{\boldsymbol{sc}}}$) introduces supervised angular margins over the unit hypersphere to enhance inter-class separability within the shared manifold, and Symmetric Cross-Entropy Loss ($\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$) jointly minimizes forward and reverse divergences between image- and prompt-based class distributions. Consequently, few-shot samples converge toward multimodal semantic equilibrium via this unified anchor-guided manifold space to facilitate class-level prompt prediction. More mathematical derivations and additional theoretical proofs are provided in the Proofs document.
@@
We use BiomedCLIP and other biomedical CLIP variants take a tokenized prompt sequence $\boldsymbol{X_{t}}\in\mathbb{R}^{B\times L}$ and a preprocessed RGB medical image $\boldsymbol{X_{v}}\in\mathbb{R}^{B\times 3\times H\times W}$ as input, and encodes them into the embedding space via a text encoder $\boldsymbol{E_{t}}$ and a vision encoder $\boldsymbol{E_{v}}$, respectively. The text $\boldsymbol{T}$ and vision embeddings $\boldsymbol{V}$ are given by $\boldsymbol{E_{t}(X_{t})}\in\mathbb{R}^{C\times D}$ and $\boldsymbol{E_{v}(X_{v})}\in\mathbb{R}^{B\times D}$, where $B$ is the batch size, $L$ is the maximum token length, $H$ and $W$ are image dimensions, $C$ is the number of classes, and $D$ is the embedding dimension. Our prompt learner constructs class‑specific prompts of the form “a photo of a[CLASS].”, where “a photo of a” serves as the initialized learnable context, and [CLASS] denotes the class name token. Classification is performed by computing cosine similarities $\cos(,)$ between the image features $V$ and each class prompt embedding $\boldsymbol{T_{p}^{(i)}}$, scaling these scores by the temperature $\tau$, applying $\mathrm{Softmax}\bigl(\cos(\boldsymbol{V},\boldsymbol{T_{p}^{(i)}})/\tau\bigr)$ to obtain class probabilities $p_{i}$, and selecting the final label as $\arg\max_{i}\,p(y=i\mid\boldsymbol{V,T_{p}^{(i)}})$. Following the setup of BiomedCoOp, we generate prompts for each class $c_{i}$ in medical modality by querying each of several recent LLMs separately with the standardized instruction:“Give N textual descriptions of visual discriminative features for distinct medical cases of [CLASS] found in [MODALITY].” We treat all LLM-provided template features as samples distributed on a shared manifold space and estimate a class-level semantic prototype field over this space. For each class $c_{i}$, we extract a semantic prototype vector $\boldsymbol{\mu}_{L,c}^{(i)}$, which guides the prompt learner in balancing multi-modal alignment and semantic expressiveness.
@@ 2. How to avoid distractors? We apply a row-wise softmax cross-entropy across $C$ classes:
Figure 2: Effect of prompting variations on saliency maps, where (a)–(e) illustrate different strategies (zoom in for details). Note that the fourth row depicts a rare cardiac cine MRI case in which the patient has a posterior mediastinal tumor, simulating a few-shot fine-tuning scenario with limited data and reflecting real clinical settings. vMFCoOp (ours) successfully localizes the approximate lesion region under such challenging conditions, while other methods, such as BiomedCoOp, tend to focus their attention on the cardiac area. This may occur because they fail to capture the concept of the posterior mediastinum when classifying “Malignant neoplasm of heart, mediastinum, and pleura”, or because their attention is overly biased toward the heart region due to semantic inductive bias, leading to overfitting or semantic misalignment.
@@
We summarize the three hierarchical constraints that define the core components of vMFCoOp. In this ablation, we systematically isolate each component under identical settings: $\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$ denotes replacing standard cross-entropy with Symmetric Cross-Entropy Loss atop BiomedCLIP; $\mathcal{L}_{\mathit{\boldsymbol{anc}}}$ introduces a unified semantic anchor field via inverse vMF estimation on the hyperspherical manifold; and $\mathcal{L}_{\mathit{\boldsymbol{sc}}}$ adds Spherical Contrastive Loss to enforce angular margins. As shown in Table 3, each constraint provides incremental improvements, while their combination yields the strongest gains. Notably, incorporating $\mathcal{L}_{\mathit{\boldsymbol{anc}}}$ consistently enhances both few-shot performance and base-to-novel generalization, underscoring the importance of aligning multimodal representations via a shared semantic anchor. Adding $\mathcal{L}_{\mathit{\boldsymbol{sc}}}$ further improves novel class accuracy, especially in higher-shot regimes, by reducing overfitting through explicit angular separation. While $\mathcal{L}_{\mathit{\boldsymbol{SCE}}}$ alone offers modest improvements, its integration with the other two constraints amplifies robustness and cross-modal alignment. These results collectively demonstrate that the three constraints are complementary and jointly contribute to optimal generalization and stability in biomedical few-shot learning. Please note that the base-to-novel results in Table 3 slightly differ from those presented in Table 2 because the BUSI dataset was excluded in Table 2, as mentioned earlier.
@@ Conclusion
Ethics Statement
The UK Biobank received ethical approval from the NHS National Research Ethics Service North West (11/NW/0382; 16/NW/0274), and no additional ethics review was required for this study.
Acknowledgments
This research was conducted using data from the UK Biobank Resource under Application Number 603483, as part of an existing approved project. The authors express their sincere gratitude to all participants and research personnel involved in the UK Biobank initiative. This work was further supported by the Scheme Towards Collaborative Cloud-Edge Deep Learning Deployment (Grant IEC/NSFC/223523), the National Edge AI Hub for Real Data: Edge Intelligence for Cyber-Disturbances and Data Quality (EP/Y028813/1), and the UK Medical Research Council (MRC) Innovation Fellowship (Grant MR/S003916/2).